# Taming Toxicity or Fueling It? The Great Ban’s Role in Shifting Toxic User Behavior and Engagement

### v1

@@ Abstract
In today’s online environments users experience harm and abuse on a daily basis. Therefore, content moderation is crucial to ensure their safety and well-being. However, the effectiveness of many moderation interventions is still uncertain. We evaluate the effectiveness of The Great Ban, one of the largest deplatforming interventions carried out by Reddit that affected almost 2,000 communities. We analyze 53M comments shared by nearly 34K users, providing in-depth results on both the intended and unintended consequences of this ban. We found that 15.6% of the moderated users abandoned the platform while the remaining ones decreased their overall toxicity by 4.1%. Nonetheless, a subset of those users increased their toxicity by 70% after the intervention. In any case, increases in toxicity did not lead to marked increases in activity or engagement, meaning that the most toxic users had overall a limited impact. Our findings bring to light new insights on the effectiveness of deplatforming. Furthermore, they also contribute to informing future content moderation strategies.
@@
content moderation , online toxicity , deplatforming , Reddit
journal: Computer Communications
@@
[2]organization=University of Pisa - Department of Computer Science,addressline=Largo Bruno Pontecorvo 2, postcode=56122, city=Pisa, country=Italy
@@ 1 Introduction
Content moderation is essential for online platforms, as it prevents the spread of harmful content and hateful behavior such as the perpetuation of hate speech [1]. It is also crucial to promote fairness and safety among users by enforcing ethical standards and supporting the health of online communities [2]. In fact, platforms use content moderation as a means to enforce their policies [3]. In particular, administrators apply different types of interventions, which can be more or less severe. For example, they can send short warning messages and use informative labels [4], or even take the drastic decision to remove content and/or users [5, 6]. However, despite the increasing reliance on content moderation, there is still little understanding of the effects of most moderation interventions, which threatens their efficacy. Recent studies have demonstrated that while some interventions had heterogeneous  [7, 8] or even ineffectual outcomes [9], others led to undesirable consequences [10, 11]. This is why it is crucial to assess the effects of recent moderation interventions as a preliminary step in planning and developing new ones.
@@
The most popular and widely used intervention is known as deplatforming, which involves the removal of content, users, or even entire communities [12]. Notorious examples are the ban that Donald Trump received in 2021 from Facebook and X (formerly Twitter) [13] and the deplatforming of three particularly toxic influencers from X  [6]. Additionally, X removed accounts involved in coordinated inauthentic behavior [14] and Reddit permanently shut down different communities because of racism, sexism and hatefulness [15, 16]. In June 2020, Reddit itself hosted one of the biggest deplatforming campaigns in the history of social media –The Great Ban– which resulted in around 2,000 subreddits being banned due to ongoing spread of toxicity and hate speech. [^1] https://www.reddit.com/r/announcements/comments/hi3oht/update_to_our_content_policy/ (accessed: 10/15/2024) Among these are popular communities such as r/The_Donald and r/ChapoTrapHouse. Despite its impact on several communities and users within and outside Reddit, its effects are still little-explored. For example, current research on The Great Ban has mainly investigated the changes in the writing style of the users [17] without assessing the ban’s effectiveness. Few studies analyzed the changes in toxicity, and those who did focused only on a small set of subreddits [5]. Moreover, the majority of the existing works on The Great Ban are focused on community-level effects, neglecting the individual user-level responses that are instrumental for understanding how effective the ban was in mitigating problematic behaviors [18, 8]. In a previous work, we provided preliminary results on the effectiveness and unintended consequences of The Great Ban [19]. Here, we extend our previous analysis by adopting a robust causal method based on Difference-in-Differences (DiD) to estimate the effects of the intervention. Furthermore, we address an additional research question investigating the impact that toxic users had after the intervention, as explained in the following.
@@
Our work provides an in-depth analysis of the effects of The Great Ban. It points out the shortcomings by highlighting the complex challenges of moderating different communities. Our findings can guide the development of future moderation interventions aimed at increasing their effectiveness.
@@ 2.1 Deplatforming
The Great Ban was a significant event that had a major impact on Reddit. However, only a few studies actually looked into its effects. One of them is the study conducted by Trujillo et al. [17], that investigated the changes in activity and use of language among the 15 most popular subreddits involved in the ban. They discovered that the most active users drastically decreased their level of activity and observed heterogeneous responses to the ban both intra- and inter- subreddits. Here, we build on this study by evaluating the impact of The Great Ban in terms of toxicity, rather than focusing on activity and language. Another body of works focused on assessing the effects of deplatforming in a subset of subreddits affected by the ban or in entirely different subreddits. Chandrasekharan et al. [15] and Saleem and Ruths [21] evaluated how The Great Ban impacted two specific subreddits: r/fatpeoplehate and r/coontown. They revealed how a large portion of users abandoned the platform and the remaining ones notably decreased their toxicity levels. However, they also observed that several users migrated to other subreddits and doubled their posting activity [22]. Other studies focused on deplatforming both within and outside Reddit. Horta Ribeiro et al. [7] studied the migration of Reddit users from banned subreddits to other platforms after a deplatforming intervention. They show how migrated users became drastically less active on the new platforms. This was not the case for a group of users, who instead became much more toxic and radicalized.  Mekacher et al. [23] analyzed the consequences of deplatforming from Twitter to Gettr. They found that politically polarized users exhibit lower toxicity on fringe platforms, likely due to reduced exposure to interactions with out-group members. Moreover, Cima et al. [14] examined large-scale bans implemented by Twitter to combat coordinated inauthentic behaviors. To conclude,  Jhaver et al. [6] studied the impact of the banishment of three notorious influencers from Twitter. They observed a general decrease in attention towards these influencers and toxicity among their fans. However, some users became much more active and toxic.
@@ 2.2 Soft moderation
Despite deplatforming being one of the most widely adopted moderation interventions, it can be perceived by many users as a threat to free speech since it involves the removal of content and users [24]. This led to the rise of an alternative category of interventions known as soft interventions. Recently, these have been the focus of several studies. For example, Trujillo and Cresci [5, 8] investigated the effects of quarantines and restrictions, which are frequently implemented before community bans on Reddit. In particular, they analyzed the effects of moderation on r/the_donald, revealing a decrease in overall activity and toxicity. However, this came at the expense of increased political polarization and decreased factual accuracy in shared news [5, 8].  Chandrasekharan et al. [22] and Shen and Rosé [25] also studied the quarantine of r/the_donald. They determined that this intervention was essentially ineffectual and did not produce notable changes in terms of misogynistic and racist comments or users’ engagement and behavior.
Another common example is flagging disputed posts with warning labels, whose effectiveness was evaluated in terms of perceived credibility and user engagement.  Pennycook et al. [11] revealed that the use of warning labels can enhance the perceived credibility of unlabeled posts, including those containing false information that has not yet been addressed. In addition,  Zannettou [24] demonstrated how users tend to reply often to tweets containing warning labels in order to further debunk the claims. However, this leads to labeled tweets circulating more than unlabelled ones. To conclude,  Katsaros et al. [4] assessed the effectiveness of showing warning messages to users before they post toxic tweets through an A/B test on Twitter. The test confirmed that this strategy is effective at reducing the number of toxic content on average, although a portion of users actually increased their toxicity.
@@ 3 Dataset
Our dataset [^2] https://doi.org/10.5281/zenodo.14034510 includes 16M comments shared by 16,828 distinct Reddit users who have been active in at least one of the 15 most popular public subreddits that were banned during The Great Ban [17], as detailed in Table 1. In this context, the term popular refers to the number of daily active users. Even though Reddit administrators initially published an obfuscated list of the most popular banned subreddits, [^3] https://www.redditstatic.com/banned-subreddits-june-2020.txt (accessed: 10/15/2024) [17] were able to decipher this list for subreddits with over 2,000 daily active users. A total of 15 subreddits were identified. The composition of our dataset is shown in Figure 1, and the procedure adopted to build it is described below.
Focus group. We began by collecting all comments posted between December 2019 and June 2020 in each of the 15 popular banned subreddits, yielding 8M comments from 194K distinct users. To this end, we used the Pushshift data archive [26]. The dataset covers 30 weeks (7 months) leading up to The Great Ban, providing a suitable baseline for user activity before the moderation intervention [8]. Notably, we were unable to collect data between May and June 2020 due to some subreddits halting their activity or being banned prior to Reddit’s public announcement of The Great Ban on June 29, 2020, as shown in Figure 1.
Following recent literature [5, 27], we identified a representative set of users for the selected subreddits by focusing on core users, that is users who consistently participated in at least one subreddit. Core users were characterized as those who posted at least one comment each month from December 2019 to March 2020.444We excluded April 2020 from this analysis due to the limited number of collected comments, likely because subreddit activity had already slowed or halted in the early days of that month. Moreover, we excluded bots (i.e., clearly automated accounts) by removing all accounts that posted two or more comments at the exact same timestamp [28]. To this end, we implemented a minimum time interval of 1 second between comments, ensuring that we did not unintentionally exclude genuine users. We verified the validity of this approach by manually validating a random sample of 1,000 comments. After these filtering steps, we ended up with 2.2M comments made by 16,828 core users. Henceforth, we denote this subset of our dataset as IN-BEFORE, representing activity within the banned subreddits prior to the ban.
@@
Baseline group. We obtain accurate causal estimates of the effects of The Great Ban by comparing the behavioral changes exhibited by the core users to possible changes exhibited during the same time by a set of reference Reddit users who were not affected by the ban. To reach this goal, we first identify a small set of related, but non-banned, subreddits to act as a baseline for the banned ones. To obtain this baseline, we first preprocessed every comment in IN-BEFORE for each banned subreddit by removing URLs, stopwords, and words less than four characters long. Subsequently, we performed stemming. For r/chapotraphouse we executed these steps only on a random sample of 1M comments, due to the sheer number of comments in the subreddit. The preprocessed comments were then fed to BERTopic  [29] to extract up to ten relevant topics, each consisting of ten keywords. We then used the 100 descriptive keywords for each subreddit to project it with BERT [30] into a 768-dimensional embedding space. This representation is suitable to highlight commonalities between the banned subreddits, such as via clustering, which we leverage to identify the small set of baseline subreddits. However, the high dimensionality of the embedding space is not suitable to directly perform clustering [31, 32]. Given the low number of data points – the 15 banned subreddits – we therefore first greatly reduce the dimensionality with PCA before clustering the banned subreddits with k-means. We experimented with multiple numbers $p$ of principal components and values for $k$, ultimately choosing $p=5$ and $k=5$ by optimizing the silhouette coefficient. Table 2 illustrates the clustering output. We now identify a set of reference non-banned subreddits for each cluster of banned subreddits, with the exception of cluster 5 that is formed by particularly small subreddits (i.e., r/imgoingtohellforthis2 and r/wojak). For clusters one to four, we computed the cluster centroid and found the most representative keywords as those having the smallest Euclidean distance to the centroid. We then retrieved all comments posted during the pre-ban period that contained those keywords and identified the non-banned subreddits where such keywords occurred more frequently  [33]. These subreddits are related to the non-banned ones and are therefore suitable candidates for representing a baseline for our causal analyses. As reported in Table 2, r/politics, r/askreddit, and r/worldnews are overall the three most representative non-banned subreddits for our clusters. As reported in Table 1, given the numerosity of the baseline subreddits, we selected a random stratified sample of 17K users, which is comparable to the number of core users in the focus group of banned subreddits. From this set, we removed 45 users who posted at least one comment in one of the 15 banned subreddits, so as to avoid any overlap between the focus and baseline groups.
@@ 4.1 RQ1: Effectiveness of The Great Ban
Figure 2 presents the aggregated results across all subreddits. Nonetheless, the same plots can be used to investigate the effects of The Great Ban among single subreddits, as done by [19]. This is useful for identifying common patterns and potential differences among the subreddits. We therefore conducted the analysis separately for users in each subreddit. The comparison across different subreddits reaffirmed earlier findings that most users who experienced significant toxicity changes after the ban increased their toxicity. However, this behavior was more pronounced in some subreddits and less evident in others. Overall, we observed that the findings for RQ1 in  [19], which utilized the median, are similar to those presented in this section when using the mean and a causal inference method like DiD.
@@ 4.2 RQ2: Extreme user reactions to The Great Ban
To further assess the behavior of outlier users – those who drastically increased or decreased their toxicity – we recompute $\Delta t^{+}$ and $\Delta t^{-}$ by only considering those users whose $|\Delta t(i)|>0.25$. Figure 4 reiterates the previous analysis for outlier users. The proportion of outlier users (left-hand side of the figure) remains relatively small in all subreddits, with r/gendercritical exhibiting the highest percentage with almost 6% of users, while r/hatecrimehoaxes does not have outliers at all. The results are notably different than those of Figure 3, as for 11 out of 15 banned subreddits (red colored) the increase in toxicity (right-hand side of the figure) significantly outweighs the decrease. This does not apply to r/gendercritical, r/soyboys, and r/oandaexlusiveforms, as for the first two subreddits, the increase in toxicity is less than 50% and no increase is observed for the latter. For what concerns the baseline subreddits, the vast majority of outliers exhibited a marked toxicity increment. However, the number of outliers within the baseline subreddits is much lower with respect to the outliers in the banned subreddits, and ranges from 0.13% to 0.22%. In addition to these results, we also repeated the analysis for multiple values of $|\Delta t(i)|$, as described in [19]. We found that the obverved behavior is independent of the way in which outlier users are defined, which supports the robusteness of our finding. Overall, our results reveal that the phenomenon of outlier users who drastically increased their toxicity post-ban is peculiar of the banned subreddits, and can possibly be explained as a negative reaction of such users to the ban itself. Moreover, our findings also underscore that extremely toxic users are spread across almost all banned subreddits, which is a sign of a systemic reaction to the ban.

### v3

# Investigating the heterogenous effects of a massive content moderation intervention via Difference-in-Differences

@@ Abstract
In today’s online environments, users encounter harm and abuse on a daily basis. Therefore, content moderation is crucial to ensure their safety and well-being. However, the effectiveness of many moderation interventions is still uncertain. Here, we apply a causal inference approach to shed light on the effectiveness of The Great Ban, a massive social media deplatforming intervention. We analyze 53M comments shared by nearly 34K users, providing in-depth results on both the intended and unintended consequences of the ban. Our causal analyses reveal that 15.6% of the moderated users abandoned the platform while the remaining ones decreased their overall toxicity by 4.1%. Nonetheless, a subset of those users increased their toxicity by 70% after the intervention. However, the increases in toxicity did not lead to marked increases in activity or engagement, meaning that the most toxic users had an overall limited impact. Our findings bring to light new insights on the effectiveness of deplatforming moderation interventions. Furthermore, they also contribute to informing future content moderation strategies.
@@
content moderation , causal inference , online toxicity , deplatforming
journal: Information Processing & Management
@@
[2]organization=University of Pisa - Department of Computer Science,addressline=Largo Bruno Pontecorvo 3, postcode=56122, city=Pisa, country=Italy
@@ 1 Introduction
Content moderation is a fundamental component of many online platforms as it prevents the spread of problematic content and behavior, such as hate speech (Giorgi et al., 2024; Kocoń et al., 2021), fake news  (Bazmi et al., 2023; Augenstein et al., 2024), and polarization  (Tao et al., 2025; Trujillo and Cresci, 2023). Content moderation also contributes to the fairness and safety of online users by promoting high ethical standards and fostering the health of online communities (Gillespie, 2018). In practice, administrators enforce platform policies through the application of a wide array of moderation interventions, which can vary in severity (Trujillo et al., 2025). For example, moderators can send short warning messages and use informative labels (Katsaros et al., 2022), or even take the drastic decision to remove content and/or users (Trujillo and Cresci, 2022; Jhaver et al., 2021; Russo et al., 2023). However, despite the increasing reliance on content moderation, there is still little understanding of the effects of many interventions, which hinders our capacity to design and apply effective moderation. For example, recent studies demonstrated that while some interventions had the desired effects, others resulted in heterogeneous (Horta Ribeiro et al., 2021; Trujillo and Cresci, 2023) or even ineffectual outcomes (Dias et al., 2020), or even led to undesirable consequences (Bail et al., 2018; Pennycook et al., 2020). Therefore, assessing the effects of moderation interventions represents an important preliminary step towards developing new and more effective forms of moderation.
@@
The most popular and widely used type of intervention is deplatforming, which involves the removal of content, users, or even entire communities (Horta Ribeiro et al., 2024). Notorious examples are the ban that Donald Trump received in 2021 from Facebook and X (formerly Twitter) (Seeliger and Baum, 2023) and the deplatforming of three particularly toxic influencers from X  (Jhaver et al., 2021). Additionally, X removed accounts involved in coordinated inauthentic behavior (Cima et al., 2024a) and Reddit permanently shut down different communities because of racism, sexism, and hatefulness (Chandrasekharan et al., 2017; Habib et al., 2022). In June 2020, Reddit itself hosted one of the biggest deplatforming campaigns in the history of social media –The Great Ban– which resulted in around 2,000 subreddits being banned due to ongoing spread of toxicity and hate speech. [^1] https://www.reddit.com/r/announcements/comments/hi3oht/update_to_our_content_policy/ (accessed: 10/15/2024) Among these are popular communities such as r/The_Donald and r/ChapoTrapHouse. Despite its impact on several communities and users within and outside Reddit, its effects are still little-explored. For example, current research on The Great Ban has mainly investigated the changes in the writing style of the users (Trujillo et al., 2021) without assessing the ban’s effectiveness. Few studies analyzed the changes in toxicity, and those who did focused only on a small set of subreddits (Trujillo and Cresci, 2022). Moreover, the majority of the existing works on The Great Ban are focused on community-level effects, neglecting the individual user-level responses that are instrumental for understanding how effective the ban was in mitigating problematic behaviors (Robertson, 2022; Trujillo and Cresci, 2023). In a previous work, we provided preliminary results on the effectiveness and unintended consequences of The Great Ban (Cima et al., 2024b). Here, we extend our previous analysis by adopting a robust causal approach based on Difference-in-Differences (DiD) to estimate the effects of the intervention. With respect to our previous work we also address an additional research question investigating the impact that toxic users had after the intervention, as explained in the following.
@@
Our work provides an in-depth analysis of the effects of The Great Ban. It points out its shortcomings by highlighting the complex challenges of moderating different communities. As such, our findings can inform the development of future moderation interventions.
@@ 2.1 Deplatforming
The Great Ban was a significant event that had a major impact on Reddit. However, only a few studies actually looked into its effects. One of them is the study conducted by Trujillo et al. (2021), that investigated the changes in activity and use of language among the 15 most popular subreddits involved in the ban. They discovered that the most active users drastically decreased their level of activity and observed heterogeneous responses to the ban both intra- and inter- subreddits. Here, we build on this study by evaluating the impact of The Great Ban in terms of toxicity, rather than focusing on activity and language. Another body of works focused on assessing the effects of deplatforming in a subset of subreddits affected by the ban or in entirely different subreddits. Chandrasekharan et al. (2017) and Saleem and Ruths (2018) evaluated how The Great Ban impacted two specific subreddits: r/fatpeoplehate and r/coontown. They revealed how a large portion of users abandoned the platform and the remaining ones notably decreased their toxicity levels. However, they also observed that several users migrated to other subreddits and doubled their posting activity (Chandrasekharan et al., 2022). Other studies focused on deplatforming both within and outside Reddit. Horta Ribeiro et al. (2021) studied the migration of Reddit users from banned subreddits to other platforms after a deplatforming intervention. They show how migrated users became drastically less active on the new platforms. This was not the case for a group of users, who instead became much more toxic and radicalized.  Mekacher et al. (2023) analyzed the consequences of deplatforming from Twitter to Gettr. They found that politically polarized users exhibit lower toxicity on fringe platforms, likely due to reduced exposure to interactions with out-group members. Moreover, Cima et al. (2024a) examined large-scale bans implemented by Twitter to combat coordinated inauthentic behaviors. To conclude,  Jhaver et al. (2021) studied the impact of the banishment of three notorious influencers from Twitter. They observed a general decrease in attention towards these influencers and toxicity among their fans. However, some users became much more active and toxic.
@@ 2.2 Soft moderation
Despite deplatforming being one of the most widely adopted moderation interventions, it can be perceived by many users as a threat to free speech since it involves the removal of content and users (Zannettou, 2021). This led to the rise of an alternative category of interventions known as soft interventions. Recently, these have been the focus of several studies. For example, Trujillo and Cresci (2022, 2023) investigated the effects of quarantines and restrictions, which are frequently implemented before community bans on Reddit. In particular, they analyzed the effects of moderation on r/the_donald, revealing a decrease in overall activity and toxicity. However, this came at the expense of increased political polarization and decreased factual accuracy in shared news (Trujillo and Cresci, 2022, 2023).  Chandrasekharan et al. (2022) and Shen and Rosé (2022) also studied the quarantine of r/the_donald. They determined that this intervention was essentially ineffectual and did not produce notable changes in terms of misogynistic and racist comments or users’ engagement and behavior.
Another common example is flagging disputed posts with warning labels, whose effectiveness was evaluated in terms of perceived credibility and user engagement.  Pennycook et al. (2020) revealed that the use of warning labels can enhance the perceived credibility of unlabeled posts, including those containing false information that has not yet been addressed. In addition,  Zannettou (2021) demonstrated how users tend to reply often to tweets containing warning labels in order to further debunk the claims. However, this leads to labeled tweets circulating more than unlabelled ones. To conclude, Katsaros et al. (2022) assessed the effectiveness of showing warning messages to users before they post toxic tweets through an A/B test on Twitter. The test confirmed that this strategy is effective at reducing the number of toxic content on average, although a small fraction of users actually increased their toxicity.
@@ 3 Dataset
Our dataset [^2] https://doi.org/10.5281/zenodo.14034510 includes 16M comments shared by 16,828 distinct Reddit users who have been active in at least one of the 15 most popular public subreddits that were banned during The Great Ban (Trujillo et al., 2021), as detailed in Table 1. In this context, the term popular refers to the number of daily active users. Even though Reddit administrators initially published an obfuscated list of the most popular banned subreddits, [^3] https://www.redditstatic.com/banned-subreddits-june-2020.txt (accessed: 10/15/2024) Trujillo et al. (2021) were able to decipher this list for subreddits with over 2,000 daily active users. A total of 15 subreddits were identified. The composition of our dataset is shown in Figure 1, and the procedure adopted to build it is described below.
Focus group. We began by collecting all comments posted between December 2019 and June 2020 in each of the 15 popular banned subreddits, yielding 8M comments from 194K distinct users. To this end, we used the Pushshift data archive (Baumgartner et al., 2020). The dataset covers 30 weeks (7 months) leading up to The Great Ban, providing a suitable baseline for user activity before the moderation intervention (Trujillo and Cresci, 2023). Notably, we were unable to collect data between May and June 2020 due to some subreddits halting their activity or being banned prior to Reddit’s public announcement of The Great Ban on June 29, 2020, as shown in Figure 1.
Following recent literature (Trujillo and Cresci, 2022; Bouleimen et al., 2023), we identified a representative set of users for the selected subreddits by focusing on core users, that is users who consistently participated in at least one subreddit. Core users were characterized as those who posted at least one comment each month from December 2019 to March 2020.444We excluded April 2020 from this analysis due to the limited number of collected comments, likely because subreddit activity had already slowed or halted in the early days of that month. Moreover, we excluded bots (i.e., clearly automated accounts) by removing all accounts that posted two or more comments at the exact same timestamp (Hurtado et al., 2019). To this end, we implemented a minimum time interval of 1 second between comments, ensuring that we did not unintentionally exclude genuine users. We verified the validity of this approach by manually validating a random sample of 1,000 comments. After these filtering steps, we ended up with 2.2M comments made by 16,828 core users. Henceforth, we denote this subset of our dataset as IN-BEFORE, representing activity within the banned subreddits prior to the ban.
@@
Baseline group. We obtain accurate causal estimates of the effects of The Great Ban by comparing the behavioral changes exhibited by the core users to possible changes exhibited during the same time by a set of reference Reddit users who were not affected by the ban. To reach this goal, we first identify a small set of related, but non-banned, subreddits to act as a baseline for the banned ones. To obtain this baseline, we first preprocessed every comment in IN-BEFORE for each banned subreddit by removing URLs, stopwords, and words less than four characters long. Subsequently, we performed stemming. For r/chapotraphouse we executed these steps only on a random sample of 1M comments, due to the sheer number of comments in the subreddit. The preprocessed comments were then fed to BERTopic  (Grootendorst, 2022) to extract up to ten relevant topics, each consisting of ten keywords. We then used the 100 descriptive keywords for each subreddit to project it with BERT (Devlin et al., 2019) into a 768-dimensional embedding space. This representation is suitable to highlight commonalities between the banned subreddits, such as via clustering, which we leverage to identify the small set of baseline subreddits. However, the high dimensionality of the embedding space is not suitable to directly perform clustering (Domingos, 2012; Zimek, 2018). Given the low number of data points – the 15 banned subreddits – we therefore first greatly reduce the dimensionality with PCA before clustering the banned subreddits with k-means (Ikotun et al., 2023). We experimented with multiple numbers $p$ of principal components and values for $k$, ultimately choosing $p=5$ and $k=5$ by optimizing the silhouette coefficient. Table 2 illustrates the clustering output. We now identify a set of reference non-banned subreddits for each cluster of banned subreddits, with the exception of cluster 5 that is formed by particularly small subreddits (i.e., r/imgoingtohellforthis2 and r/wojak). For clusters one to four, we computed the cluster centroid and found the most representative keywords as those having the smallest Euclidean distance to the centroid. We then retrieved all comments posted during the pre-ban period that contained those keywords and identified the non-banned subreddits where such keywords occurred more frequently. These subreddits are related to the non-banned ones and are therefore suitable candidates for representing a baseline for our causal analyses. As reported in Table 2, r/politics, r/askreddit, and r/worldnews are overall the three most representative non-banned subreddits for our clusters. As reported in Table 1, given the numerosity of the baseline subreddits, we selected a random stratified sample of 17K users, which is comparable to the number of core users in the focus group of banned subreddits. From this set, we removed 45 users who posted at least one comment in one of the 15 banned subreddits, so as to avoid any overlap between the focus and baseline groups.
@@ 4.1 RQ1: Effectiveness of The Great Ban
Figure 2 presents the aggregated results across all subreddits. Nonetheless, the same plots can be used to investigate the effects of The Great Ban among single subreddits, as done by Cima et al. (2024b). This is useful for identifying common patterns and potential differences among the subreddits. We therefore conducted the analysis separately for users in each subreddit. The comparison across different subreddits reaffirmed earlier findings that most users who experienced significant toxicity changes after the ban increased their toxicity. However, this behavior was more pronounced in some subreddits and less evident in others. Overall, we observed that the findings for RQ1 in  (Cima et al., 2024b), which utilized the median, are similar to those presented in this section when using the mean and a causal inference method like DiD.
@@ 4.2 RQ2: Extreme user reactions to The Great Ban
To further assess the behavior of outlier users – those who drastically increased or decreased their toxicity – we recompute $\Delta t^{+}$ and $\Delta t^{-}$ by only considering those users whose $|\Delta t(i)|>0.25$. Figure 4 reiterates the previous analysis for outlier users. The proportion of outlier users (left-hand side of the figure) remains relatively small in all subreddits, with r/gendercritical exhibiting the highest percentage with almost 6% of users, while r/hatecrimehoaxes does not have outliers at all. The results are notably different than those of Figure 3, as for 11 out of 15 banned subreddits (red colored) the increase in toxicity (right-hand side of the figure) significantly outweighs the decrease. This does not apply to r/gendercritical, r/soyboys, and r/oandaexlusiveforms, as for the first two subreddits, the increase in toxicity is less than 50% and no increase is observed for the latter. For what concerns the baseline subreddits, the vast majority of outliers exhibited a marked toxicity increment. However, the number of outliers within the baseline subreddits is much lower with respect to the outliers in the banned subreddits, and ranges from 0.13% to 0.22%. In addition to these results, we also repeated the analysis for multiple values of $|\Delta t(i)|$, as described in Cima et al. (2024b). We found that the obverved behavior is independent of the way in which outlier users are defined, which supports the robusteness of our finding. Overall, our results reveal that the phenomenon of outlier users who drastically increased their toxicity post-ban is peculiar of the banned subreddits, and can possibly be explained as a negative reaction of such users to the ban itself. Moreover, our findings also underscore that extremely toxic users are spread across almost all banned subreddits, which is a sign of a systemic reaction to the ban.